In 2012, the imaginary nation of Bovina had a population of 8,044 and real GDP of 36,198,000. In 2013 it had a population of 7,800 and real GDP of 35,880,000. What is the growth rate of real GDP per person in Bovina between 2013 and 1012?
Answer #1

Ans: 2.22%

Explanation:

Real GDP per capita = Real GDP / Total population of the country

Real GDP per capiat in 2102 = 36,198,000 / 8,044 = 4500

Real GDP per capita in 2103  = 35,880,000/ 7800 = 4600

The growth rate of real GDP per person in Bovina between 2013 and 1012;

= {( 4600 -4500 ) / 4500 } *100

= ( 100 / 4500 ) *100

= 0.0222 *100

= 2.22%